# NeAR: Coupled Neural Asset–Renderer Stack

### v1

@@ 3.3 LH-SLAT Reconstruction & Generation
ADT [32] DTC [9] Objaverse data [8] Glossy Synthetic dataset [24]
@@
3.4 Relightable Neural 3DGS Synthesis
The second stage synthesizes a relightable 3D Gaussian Splatting (3DGS) field $\mathcal{G}$ from LH-SLAT, conditioned on target illumination and viewpoint. Unlike iterative optimization approaches [12, 2], we employ an efficient feed-forward decoder with two sequential modules: the Intrinsic Aware Decoder (IAD) and the Lighting Aware Decoder (LAD).
@@ 3.4.1 Intrinsic Aware Decoder (IAD)
The IAD, denoted as $\mathcal{D}_{I}$, aims to process LH-SLAT $Z_{\text{lh}}$ and generate a view-independent and illumination-invariant intrinsic feature $\bm{h}=\{(\bm{h}_{i},\bm{p}_{i})\}_{i=1}^{L}$, where $\bm{h}_{i}\in\mathbb{R}^{768}$. This sparse feature field $\bm{h}$ effectively decodes the underlying geometric structure and material properties of the scene. To achieve this, IAD employs a Transformer architecture akin to TRELLIS [45], leveraging stacked self-shifted window attention blocks to exploit the inherent locality of structured 3D latent sequences. To further enhance the model’s comprehension of global structural relationships and lighting context, a register cross-attention layer is incorporated into each block. Specifically, 16 learnable register tokens are appended to each object’s corresponding SLAT token sequences. These tokens encode global scene information and potentially attenuate high-frequency noise within the embeddings [7, 19]. Finally, these register tokens are injected into the decoder via a global cross-attention mechanism, facilitating information exchange between the register tokens and all latent variable tokens, thereby enabling the generation of a coherent and globally consistent intrinsic representation $\bm{h}$.
@@ 3.4.2 Lighting Aware Decoder (LAD)
The LAD, denoted as $\mathcal{D}_{E}$, synthesizes the final lighting-dependent features by injecting view embeddings and environmental lighting conditions, as showed in Fig. 3. .
@@
Lighting Tokenizer. We encode the high dynamic range (HDR) environment map $\mathbf{E}$ into compact lighting conditions using an HDRI encoder $\mathcal{E}_{l}$. Following [16, 23, 13], we decompose $\mathbf{E}$ into a tone-mapped LDR image $\mathbf{E}_{\text{ldr}}$, a normalized log-intensity map $\mathbf{E}_{\text{log}}=\log(\mathbf{E}+1)/\mathbf{E}_{max}$, and a camera-space direction encoding $\mathbf{E}_{\text{dir}}\in\mathbb{R}^{H\times W\times 3}$. Unlike prior works that compress the entire map via VAE, we employ a ConvNeXt backbone to extract multi-scale visual features from $\mathbf{E}_{\text{ldr}}$ and $\mathbf{E}_{\text{log}}$. Crucially, rather than directly compressing $\mathbf{E}_{\text{dir}}$, we first encode it via NeRF-style positional embedding [30] and fuse it with visual features using Spatial Cross Attention. This mechanism acts as a learnable positional encoding, modulating visual features with explicit directional cues. The resulting multi-scale features are concatenated, processed with positional encoding, and passed through self-attention blocks to yield the Lighting Condition Tokens $C_{L}\in\mathbb{R}^{4096\times 768}$. This design explicitly embeds directional information, facilitating editable lighting directions when switching views.
@@
3D Gaussian Decoding. We regress the 3DGS parameters using both the intrinsic feature $\bm{h}$ and the lighting-aware feature $\bm{h}^{e}$:
@@
the intrinsic feature $\bm{h}_{i}$ is decoded into $K$ Gaussian parameters: position offset $\bm{o}$, base color $\bm{b}$, roughness $\gamma$, metallic $\bm{m}$, scale $\bm{s}$, rotation $\bm{r}$, and opacity $\alpha$ (activated via $\tanh$ to support negative density [61]). Simultaneously, the lighting-dependent feature $\bm{h}^{e}_{i}$ predicts the 48-dim color feature $\bm{f}$, lighting-specific scale $\hat{s}$, and shadow $\sigma$. The Gaussian centers are defined as $\bm{x}^{k}_{i}=\bm{p}_{i}+\tanh(\bm{o}^{k}_{i})$, with normals derived from the shortest axis of $\hat{s}$. Finally, we employ a simple shallow MLP network that combines the positional encoding of the normal vector and the color feature $\bm{f}$. This network uses ReLU activation functions in its intermediate layers and an ELU activation function in its final layer to predict the radiance values for each Gaussian. Through the rasterization operation $\mathcal{M}$, we obtain the 2D HDR prediction $I^{hdr}_{target}$. We also render 2D base color, roughness, metallic, shadow images $I^{b},I^{r},I^{m},I^{s}$.
Loss Function. We supervise the training via an HDR reconstruction loss $\mathcal{L}_{hdr}$, which comprises $\mathcal{L}_{1}$, LPIPS [57], D-SSIM and regularization terms. Following [53], to prevent high-intensity regions from dominating the $\mathcal{L}_{1}$ optimization, we apply a logarithmic transformation to the HDR images. For perceptual metrics (LPIPS and D-SSIM), we operate on tone-mapped images using $\text{clamp}(\log_{2}(I),0,1)$. Additionally, we impose auxiliary $\mathcal{L}_{1}$ supervision on material properties maps (base color, roughness, metallic), denoted as $\mathcal{L}_{pbr}$, and shadows $\mathcal{L}_{shadow}$. The total objective is formulated as follows:
@@ 4.1 Implementation Details
Training Data. Our training dataset comprises 87K 3D assets with physically-based rendering (PBR) textures, curated from the Objaverse-XL dataset. These assets are illuminated using 2K High Dynamic Range Images (HDRIs), each at 4K resolution, used as environment maps. We normalized the assets to fit within a bounding box of [0.5,0.5][-0.5,0.5]. The first training stage involves rendering 150 viewpoints under normalized lighting to extract illumination-invariant structural latent representations. For input images under unknown illumination, camera poses are sampled with yaw within ±45 degrees and pitch from -10 to 45 degrees, oriented towards the object’s center, and with field of view (FOV) and radius following [45]. Unknown illumination is modeled with (1) six area lights uniformly distributed on a sphere, (2) 1-3 area lights randomly sampled within the camera’s hemisphere, or (3) a random, Z-axis-rotated environment map. Area light intensities are sampled uniformly between 300 and 700 (units), distances between 5 and 8 units. In the second stage, we re-light objects using randomly rotated environment maps as supervision, with a fixed FOV of $40^{\circ}$. We randomly and uniformly sample 12 camera viewpoints on a sphere of radius 2.0, where each viewpoint is rendered under 16 different illumination conditions. All data generation across both stages utilizes the Blender EEVEE Next engine [42] with raytracing enabled.
@@
Evaluation Metric. We use PSNR, SSIM [43] and LPIPS [57] to measure the quality of the rendering.
Evaluation Datasets. We construct a test set by randomly selecting 800 unseen objects from our training data. To validate generalization capability, we evaluate on out-of-domain datasets: Aria Digital Twin (ADT) [32] and Digital Twin Catalog (DTC) [9], which feature high-fidelity photorealistic models with sub-millimeter accuracy. We also incorporate the Glossy Synthetic dataset [24] and additional assets from BlenderKit [^1] https://www.blenderkit.com/, modifying rendering nodes to utilize the Principled BSDF shader [^2] https://www.blender.org/.
@@ 4.2 Single-view Forward Rendering
Variants of $\mathcal{D}$. Tab. 2 indicates that increasing the depth of $\mathcal{D}_{E}$ improves quality but reduces inference speed; we therefore select 6 layers to strike a balance between efficiency and performance. Relying solely on the LAD $\mathcal{D}_{E}$ leads to a significant decline in relighting performance, consistent with [53]. Furthermore, Fig. 8 demonstrates that injecting camera view information to identify which lighting tokens should be attended to, prior to lighting baking, significantly enhances relighting results compared to baking global lighting first. This design allows for more effective capture of geometric and lighting variations, boosting the performance of $\mathcal{D}_{E}$ (Tab. 4).
@@ A.1 Implementation Details
In the LH-SLAT Reconstruction & Generation phase3.3), we fine-tune a rectified flow model equipped with LoRA [15] to normalize shaded SLATs from arbitrary images into LH-SLATs. The goal of this phase is to learn a mapping, $f_{\theta}$, that transforms light-dependent shaded SLAT representations into light-homogenized counterparts. To achieve this efficiently while preserving the prior knowledge of the original flow model $f_{s}$ [45], we initialize LoRA using PEFT [29]. We configure LoRA with a rank of 512 and a scaling factor, further integrating rslora [17] to enhance training stability. LoRA adaptors are applied to the query, key, value, and output projection modules within the attention mechanism. We optimize the model using AdamW [27] with a learning rate of $1\times 10^{-4}$. This training phase takes approximately two days to complete.
In the Relightable Neural 3DGS Synthesis phase3.4), we employ the AdamW optimizer with a batch size of 48. The learning rate is warmed up linearly to $1\times 10^{-4}$ over the first 5K steps, followed by a cosine decay schedule. We perform end-to-end joint training on the IAD, LAD, and the Lighting-Aware tokenizer (denoted as $E_{l}$). To accelerate training, we leverage Flash-Attention 3 [37] and gsplat [49]. The model is trained for 500K iterations across all loss components, requiring approximately 10 days. Additionally, we investigated the incorporation of geometric constraint losses, specifically normal and depth losses. However, we observed that adding these regularizations degraded both convergence speed and rendering quality, suggesting a trade-off between geometric constraints and rendering fidelity within the 3DGS framework [4, 50].
@@
Given a single input image $I_{in}$ with unknown lighting and a target high dynamic range (HDR) environment map $E$, our inference pipeline proceeds as follows. Since our method decouples geometry generation from relighting, we first reconstruct a 3D mesh $m$ from $I_{in}$ using Hunyuan3D 2.1 (Hy3D 2.1) [60] with default settings. This mesh is then voxelized to provide coordinates for the structurally sparse voxel feature SLAT.
Following Trellis [45], we utilize the pre-trained SLAT flow model $f_{s}$ to generate an initial shaded SLAT $Z_{s}$ from $I_{in}$. Note that $Z_{s}$ inherently contains arbitrary lighting information from the input image. To remove these lighting effects, we concatenate $Z_{s}$ with noise (matching $Z_{s}$ in shape) along the channel dimension and feed the result into our fine-tuned corrective model $f_{\theta}$ to yield the Lighting-Homogenized SLAT (LH-SLAT).
Subsequently, for the target lighting, we pre-process the environment map $E$ (as detailed in Sec. 3.4.2) and encode it into a lighting condition embedding $C_{L}$ using the Lighting-Aware tokenizer $\mathcal{E}_{l}$. The IAD module then processes the LH-SLAT to extract intrinsic features $h$. Simultaneously, the LAD $\mathcal{D}_{E}$ integrates the viewing direction encoding $e^{v}$ and the lighting condition $C_{L}$ to predict the 3D Gaussian attributes for the specific view and lighting (Eq. 5). Finally, we render the relit HDR image $I^{hdr}_{target}$ via gsplat [49]. To align the visual output with standard rendering engines like Blender, we apply AgX tone mapping [^3] https://github.com/iamNCJ/simple-ocio to convert the HDR result into a low dynamic range (LDR) image.
@@ A.2 Network Architectures
Apart from the lighting-aware tokenizer $\mathcal{E}_{l}$, and consistent with [45], our method primarily employs Transformer networks. As depicted in Fig. 3, the IAD $\mathcal{D}_{I}$ comprises 3D shifted window multi-head self-attention (3D-SW-MSA) and a feed-forward network (FFN). Addressing the limitation of the naive 3D-SW-MSA design in [45], which computes attention solely within local windows and neglects inter-window information exchange, we introduce learnable register tokens. These tokens interact with all windows via 3D multi-head cross-attention (3D-MCA), serving as a global information bridge to facilitate the model’s learning of global context. The lighting-aware decoder $\mathcal{D}_{E}$ receives intrinsic features $h$, view encoding $e^{v}$, register tokens, and lighting encoding to generate lighting-dependent features $h^{v}$. Register tokens and lighting encoding are injected into the network via 3D-MCA. Here, $h$ and $e^{v}$ are added in a voxel-wise manner to determine which lighting encoding tokens should be attended to under the current viewpoint. The ablation study on the interaction order of viewpoint and lighting information is illustrated in Fig. 8 and Tab 4.
@@
Regularization. To prevent the degeneration of Gaussian primitives (e.g., becoming too large or too opaque) during optimization [45], we incorporate a volumetric loss $\mathcal{L}_{\text{vol}}$ and an opacity loss $\mathcal{L}_{\alpha}$:
@@
The lighting-aware tokenizer, $\mathcal{E}_{l}$, is primarily designed to process and inject lighting information into the network to enable relighting while also effectively perceiving rotations in the environment map. Similar to Neural Graffer and Diffusion Renderer, as depicted in Fig. 9, our approach leverages the $E_{hdr}$ and $E_{log}$ components of environmental illumination to provide lighting color features. Neural Graffer encodes the environment map into the image latent space using a pretrained image VAE’s encoder. Diffusion Renderer, in contrast, employs a video VAE model to compress $E_{ldr}$, $E_{log}$, and $E_{dir}$ into a video latent space of consecutive frames, thus accommodating subsequent image or video diffusion model training. As shown in Fig. 9(c), the design of $\mathcal{E}_{l}$ aims to facilitate the injection of lighting information from the LH-SLAT into relightable 3D Gaussian Splatting (GS). This design addresses two key challenges: First, different materials require sensitivity to varying resolutions of lighting information. For instance, highly rough surfaces require only low-resolution environment maps, whereas high-metallicity surfaces with low roughness necessitate high-resolution maps. To address this, we utilize ConvNext [25] to extract multi-resolution features from the lighting pyramid and employ a spatial attention mechanism to compute attention scores and exchange information between these resolutions. Second, the model should accurately perceive rotations of the environment map. Neural Graffer requires deforming the environment map itself. Our approach, similar to Diffusion Renderer, can rotate the illumination by adjusting the environment light direction map, $E_{dir}$, requiring only the application of a rotation matrix to the direction vector. However, Diffusion Renderer relies on an additionally trained environment encoder (Env. Encoder). Our method employs a direction-encoding-aware spatial cross-multihead attention (Spatial Cross MHA) to guide visual features at different resolutions using directional information. It combines multi-scale feature fusion to preserve both detailed and global information and utilizes a RoPE+RMSNorm transformer layer for efficient sequence modeling, refer to Fig. 3. This allows the complex HDRI lighting information to be encoded into conditional tokens suitable for cross-attention, providing high-quality lighting conditions for the renderer. Abstractly, we model the environment map as a set of light source tokens, each encoded with absolute direction vector positional information and multi-scale features. Subsequently, the Lighting-Aware Decoder (LAD), $\mathcal{E}_{l}$, can efficiently determine the relevance of each token to the current viewpoint by leveraging viewpoint direction encoding.

### v2

@@ 3.4 LH-SLAT Generation
ADT [pan2023aria] DTC [dong2025digital] Objaverse data [deitke2023objaverse] Glossy Synthetic dataset [nero]
@@
Figure 5: The network structures for Lighting Tokenizer, IAD and LAD.
3.5 Relightable Neural 3D Gaussian Synthesis
The second stage synthesizes a relightable 3D Gaussian Splatting (3DGS) field $\mathcal{G}$ from LH-SLAT, conditioned on target illumination and viewpoint. Unlike iterative optimization approaches [gao2024relightable, bi2024gs3], we employ an efficient feed-forward decoder with two sequential modules: the Intrinsic Aware Decoder (IAD) and the Lighting Aware Decoder (LAD).
@@ 3.5.1 Intrinsic Aware Decoder (IAD)
The IAD, denoted as $\mathcal{D}_{I}$, aims to process LH-SLAT $Z_{\text{lh}}$ and generate a view-independent and illumination-invariant intrinsic feature $\bm{h}=\{(\bm{h}_{i},\bm{p}_{i})\}_{i=1}^{L}$, where $\bm{h}_{i}\in\mathbb{R}^{768}$. This sparse feature field $\bm{h}$ effectively decodes the underlying geometric structure and material properties of the scene. To achieve this, IAD employs a Transformer architecture akin to TRELLIS [xiang2024structured], leveraging stacked self-shifted window attention blocks to exploit the inherent locality of structured 3D latent sequences. To further enhance the model’s comprehension of global structural relationships and lighting context, a register cross-attention layer is incorporated into each block. Specifically, 16 learnable register tokens are appended to each object’s corresponding SLAT token sequences. These tokens encode global scene information and potentially attenuate high-frequency noise within the embeddings [darcet2024vision, li2025lino]. Finally, these register tokens are injected into the decoder via a global cross-attention mechanism, facilitating information exchange between the register tokens and all latent variable tokens, thereby enabling the generation of a coherent and globally consistent intrinsic representation $\bm{h}$.
@@ 3.5.2 Lighting Aware Decoder (LAD)
The LAD, denoted as $\mathcal{D}_{E}$, synthesizes the final lighting-dependent features by injecting view embeddings and environmental lighting conditions, as shown in Fig. 4.
@@
Lighting Tokenizer. We encode the high dynamic range (HDR) environment map $\mathbf{E}$ into compact lighting conditions using an HDRI encoder $\mathcal{E}_{l}$. Following [jin2024neural, diffusionrenderer, unirelight], we decompose $\mathbf{E}$ into a tone-mapped LDR image $\mathbf{E}_{\text{ldr}}$, a normalized log-intensity map $\mathbf{E}_{\text{log}}=\log(\mathbf{E}+1)/\mathbf{E}_{max}$, and a camera-space direction encoding $\mathbf{E}_{\text{dir}}\in\mathbb{R}^{H\times W\times 3}$. Unlike prior works that compress the entire map via VAE, we employ a ConvNeXt backbone to extract multi-scale visual features from $\mathbf{E}_{\text{ldr}}$ and $\mathbf{E}_{\text{log}}$. Crucially, rather than directly compressing $\mathbf{E}_{\text{dir}}$, we first encode it via NeRF-style positional embedding [mildenhall2021nerf] and fuse it with visual features using Spatial Cross Attention. This mechanism acts as a learnable positional encoding, modulating visual features with explicit directional cues. The resulting multi-scale features are concatenated, processed with positional encoding, and passed through self-attention blocks to yield the Lighting Condition Tokens $C_{L}\in\mathbb{R}^{4096\times 768}$. This design explicitly embeds directional information, facilitating editable lighting directions when switching views.
@@
3.5.3 Neural 3D Gaussian Splatting
We regress the 3DGS parameters using both the intrinsic feature $\bm{h}$ and the lighting-aware feature $\bm{h}^{e}$:
@@ 3.5.3 Neural 3D Gaussian Splatting
the intrinsic feature $\bm{h}_{i}$ is decoded into $K$ Gaussian parameters: position offset $\bm{o}$, base color $\bm{b}$, roughness $\gamma$, metallic $\bm{m}$, scale $\bm{s}$, rotation $\bm{r}$, and opacity $\alpha$ (activated via $\tanh$ to support negative density [zhu20253d]). Simultaneously, the lighting-dependent feature $\bm{h}^{e}_{i}$ predicts the 48-dim color feature $\bm{f}$, lighting-specific scale $\hat{s}$, and shadow $\sigma$. The Gaussian centers are defined as $\bm{x}^{k}_{i}=\bm{p}_{i}+\tanh(\bm{o}^{k}_{i})$, with normals derived from the shortest axis of $\hat{s}$. Finally, we employ a simple shallow MLP network that combines the positional encoding of the normal vector and the color feature $\bm{f}$. This network uses ReLU activation functions in its intermediate layers and an ELU activation function in its final layer to predict the radiance values for each Gaussian. Through the rasterization operation $\mathcal{M}$, we obtain the 2D HDR prediction $I^{hdr}_{target}$. We also render 2D base color, roughness, metallic, shadow images $I^{b},I^{r},I^{m},I^{s}$.
Loss Function. We supervise the training via an HDR reconstruction loss $\mathcal{L}_{hdr}$, which comprises $\mathcal{L}_{1}$, LPIPS [zhang2018perceptual], D-SSIM and regularization terms. Following [zeng2025renderformer], to prevent high-intensity regions from dominating the $\mathcal{L}_{1}$ optimization, we apply a logarithmic transformation to the HDR images. For perceptual metrics (LPIPS and D-SSIM), we operate on tone-mapped images using $\text{clamp}(\log_{2}(I),0,1)$. Additionally, we impose auxiliary $\mathcal{L}_{1}$ supervision on material properties maps (base color, roughness, metallic), denoted as $\mathcal{L}_{pbr}$, and shadows $\mathcal{L}_{shadow}$. The total objective is formulated as follows:
@@ 4.1 Implementation Details
Figure 6: Schematic illustration of four distinct sub-tasks.
Training Data. Our training dataset comprises 87K 3D assets with physically-based rendering (PBR) textures, curated from the Objaverse-XL dataset. These assets are illuminated using 2K High Dynamic Range Images (HDRIs), each at 4K resolution, used as environment maps. We normalized the assets to fit within a bounding box of [0.5,0.5][-0.5,0.5]. The first training stage involves rendering 150 viewpoints under normalized lighting to extract illumination-invariant structural latent representations. For input images under unknown illumination, camera poses are sampled with yaw within ±45 degrees and pitch from -10 to 45 degrees, oriented towards the object’s center, and with field of view (FOV) and radius following [xiang2024structured]. Unknown illumination is modeled with (1) six area lights uniformly distributed on a sphere, (2) 1-3 area lights randomly sampled within the camera’s hemisphere, or (3) a random, Z-axis-rotated environment map. Area light intensities are sampled uniformly between 300 and 700 (units), distances between 5 and 8 units. In the second stage, we re-light objects using randomly rotated environment maps as supervision, with a fixed FOV of $40^{\circ}$. We randomly and uniformly sample 12 camera viewpoints on a sphere of radius 2.0, where each viewpoint is rendered under 16 different illumination conditions. All data generation across both stages utilizes the Blender EEVEE Next engine [eevee] with raytracing enabled.
@@
Evaluation Metric. We use PSNR, SSIM [wang2004image] and LPIPS [zhang2018perceptual] to measure the quality of the rendering.
Evaluation Datasets. We construct a test set by randomly selecting 800 unseen objects from our training data. To validate generalization capability, we evaluate on out-of-domain datasets: Aria Digital Twin (ADT) [pan2023aria] and Digital Twin Catalog (DTC) [dong2025digital], which feature high-fidelity photorealistic models with sub-millimeter accuracy. We also incorporate the Glossy Synthetic dataset [nero] and additional assets from BlenderKit [^1] https://www.blenderkit.com/, modifying rendering nodes to utilize the Principled BSDF shader [^2] https://www.blender.org/.
@@ 4.2 Single-view Forward Rendering
Variants of $\mathcal{D}$. Tab. 2 indicates that increasing the depth of $\mathcal{D}_{E}$ improves quality but reduces inference speed; we therefore select 6 layers to strike a balance between efficiency and performance. Relying solely on the LAD $\mathcal{D}_{E}$ leads to a significant decline in relighting performance, consistent with [zeng2025renderformer]. Furthermore, Fig. 11 demonstrates that injecting camera view information to identify which lighting tokens should be attended to, prior to lighting baking, significantly enhances relighting results compared to baking global lighting first. This design allows for more effective capture of geometric and lighting variations, boosting the performance of $\mathcal{D}_{E}$ (Tab. 4).
@@ A.1 Implementation Details
In the LH-SLAT Reconstruction & Generation phase3.3), we fine-tune a rectified flow model equipped with LoRA [hu2022lora] to normalize shaded SLATs from arbitrary images into LH-SLATs. The goal of this phase is to learn a mapping, $f_{\theta}$, that transforms light-dependent shaded SLAT representations into light-homogenized counterparts. To achieve this efficiently while preserving the prior knowledge of the original flow model $f_{s}$ [xiang2024structured], we initialize LoRA using PEFT [peft]. We configure LoRA with a rank of 512 and a scaling factor, further integrating rslora [rslora] to enhance training stability. LoRA adaptors are applied to the query, key, value, and output projection modules within the attention mechanism. We optimize the model using AdamW [loshchilov2018decoupled] with a learning rate of $1\times 10^{-4}$. This training phase takes approximately two days to complete.
In the Relightable Neural 3DGS Synthesis phase3.5), we employ the AdamW optimizer with a batch size of 48. The learning rate is warmed up linearly to $1\times 10^{-4}$ over the first 5K steps, followed by a cosine decay schedule. We perform end-to-end joint training on the IAD, LAD, and the Lighting-Aware tokenizer (denoted as $E_{l}$). To accelerate training, we leverage Flash-Attention 3 [flash3] and gsplat [ye2025gsplat]. The model is trained for 500K iterations across all loss components, requiring approximately 10 days. Additionally, we investigated the incorporation of geometric constraint losses, specifically normal and depth losses. However, we observed that adding these regularizations degraded both convergence speed and rendering quality, suggesting a trade-off between geometric constraints and rendering fidelity within the 3DGS framework [pgsr, gof].
@@
Given a single input image $I_{in}$ with unknown lighting and a target high dynamic range (HDR) environment map $E$, our inference pipeline proceeds as follows. Since our method decouples geometry generation from relighting, we first reconstruct a 3D mesh $m$ from $I_{in}$ using Hunyuan3D 2.1 (HY3D 2.1) [zhao2025hunyuan3d] with default settings. This mesh is then voxelized to provide coordinates for the structurally sparse voxel feature SLAT.
Following Trellis [xiang2024structured], we utilize the pre-trained SLAT flow model $f_{s}$ to generate an initial shaded SLAT $Z_{s}$ from $I_{in}$. Note that $Z_{s}$ inherently contains arbitrary lighting information from the input image. To remove these lighting effects, we concatenate $Z_{s}$ with noise (matching $Z_{s}$ in shape) along the channel dimension and feed the result into our fine-tuned corrective model $f_{\theta}$ to yield the Lighting-Homogenized SLAT (LH-SLAT).
Subsequently, for the target lighting, we pre-process the environment map $E$ (as detailed in Sec. 3.5.2) and encode it into a lighting condition embedding $C_{L}$ using the Lighting-Aware tokenizer $\mathcal{E}_{l}$. The IAD module then processes the LH-SLAT to extract intrinsic features $h$. Simultaneously, the LAD $\mathcal{D}_{E}$ integrates the viewing direction encoding $e^{v}$ and the lighting condition $C_{L}$ to predict the 3D Gaussian attributes for the specific view and lighting (Eq. 5). Finally, we render the relit HDR image $I^{hdr}_{target}$ via gsplat [ye2025gsplat]. To align the visual output with standard rendering engines like Blender, we apply AgX tone mapping [^3] https://github.com/iamNCJ/simple-ocio to convert the HDR result into a low dynamic range (LDR) image.
@@ A.2 Network Architectures
Apart from the lighting-aware tokenizer $\mathcal{E}_{l}$, and consistent with [xiang2024structured], our method primarily employs Transformer networks. As depicted in Fig. 4, the IAD $\mathcal{D}_{I}$ comprises 3D shifted window multi-head self-attention (3D-SW-MSA) and a feed-forward network (FFN). Addressing the limitation of the naive 3D-SW-MSA design in [xiang2024structured], which computes attention solely within local windows and neglects inter-window information exchange, we introduce learnable register tokens. These tokens interact with all windows via 3D multi-head cross-attention (3D-MCA), serving as a global information bridge to facilitate the model’s learning of global context. The lighting-aware decoder $\mathcal{D}_{E}$ receives intrinsic features $h$, view encoding $e^{v}$, register tokens, and lighting encoding to generate lighting-dependent features $h^{v}$. Register tokens and lighting encoding are injected into the network via 3D-MCA. Here, $h$ and $e^{v}$ are added in a voxel-wise manner to determine which lighting encoding tokens should be attended to under the current viewpoint. The ablation study on the interaction order of viewpoint and lighting information is illustrated in Fig. 11 and Tab 4.
@@
Regularization. To prevent the degeneration of Gaussian primitives (e.g., becoming too large or too opaque) during optimization [xiang2024structured], we incorporate a volumetric loss $\mathcal{L}_{\text{vol}}$ and an opacity loss $\mathcal{L}_{\alpha}$:
@@
The lighting-aware tokenizer, $\mathcal{E}_{l}$, is primarily designed to process and inject lighting information into the network to enable relighting while also effectively perceiving rotations in the environment map. Similar to Neural Graffer and Diffusion Renderer, as depicted in Fig. 12, our approach leverages the $E_{hdr}$ and $E_{log}$ components of environmental illumination to provide lighting color features. Neural Graffer encodes the environment map into the image latent space using a pretrained image VAE’s encoder. Diffusion Renderer, in contrast, employs a video VAE model to compress $E_{ldr}$, $E_{log}$, and $E_{dir}$ into a video latent space of consecutive frames, thus accommodating subsequent image or video diffusion model training. As shown in Fig. 12(c), the design of $\mathcal{E}_{l}$ aims to facilitate the injection of lighting information from the LH-SLAT into relightable 3D Gaussian Splatting (GS). This design addresses two key challenges: First, different materials require sensitivity to varying resolutions of lighting information. For instance, highly rough surfaces require only low-resolution environment maps, whereas high-metallicity surfaces with low roughness necessitate high-resolution maps. To address this, we utilize ConvNext [liu2022convnet] to extract multi-resolution features from the lighting pyramid and employ a spatial attention mechanism to compute attention scores and exchange information between these resolutions. Second, the model should accurately perceive rotations of the environment map. Neural Graffer requires deforming the environment map itself. Our approach, similar to Diffusion Renderer, can rotate the illumination by adjusting the environment light direction map, $E_{dir}$, requiring only the application of a rotation matrix to the direction vector. However, Diffusion Renderer relies on an additionally trained environment encoder (Env. Encoder). Our method employs a direction-encoding-aware spatial cross-multihead attention (Spatial Cross MHA) to guide visual features at different resolutions using directional information. It combines multi-scale feature fusion to preserve both detailed and global information and utilizes a RoPE+RMSNorm transformer layer for efficient sequence modeling, refer to Fig. 4. This allows the complex HDRI lighting information to be encoded into conditional tokens suitable for cross-attention, providing high-quality lighting conditions for the renderer. Abstractly, we model the environment map as a set of light source tokens, each encoded with absolute direction vector positional information and multi-scale features. Subsequently, the Lighting-Aware Decoder (LAD), $\mathcal{E}_{l}$, can efficiently determine the relevance of each token to the current viewpoint by leveraging viewpoint direction encoding.